xXMarcelieXx:

Help please with vectors Find | a - b | a= < -3 , 4 > , b = < 9, -1>

moreso:

Just to clarify what sillybilly said $$ \| \mathbf a- \mathbf b \| ^2 = (\mathbf a - \mathbf b ) .(\mathbf a - \mathbf b ) $$Therefore $$ \| \mathbf a- \mathbf b \| = \sqrt{ (\mathbf a - \mathbf b ) .(\mathbf a - \mathbf b ) } $$

moreso:

The latter can be written more condensely as $$\| \mathbf a- \mathbf b \| = \sqrt{ (\mathbf a - \mathbf b )^2 }$$ with the understanding that multiplication here is scalar dot product

Nnesha:

\[\large\rm (a-b) \cdot (a-b)\] `\cdot`

Nnesha:

why did you delete his work ?? -_-